A pendulum clock loses 13.80 s every day. What fractional change ΔL/L must be made to the length L of the pendulum so that the clock keeps perfect time?
Answer #1

The fractional change is

ΔL/L=2(ΔT/T)

       =2(13.8)/3600*24=3.2*10-4